# Continuum-Field-Theoretical Construction of Conserved Currents, Non-Invertible Symmetries, and Mixed Anomalies in (3+1)D Non-Abelian Topological Order

### v2

@@ IV.3 An overview of fusion rules of symmetry operators
A general invertible symmetry operator is denoted by $U_{e_{1}e_{2}e_{3}}\left(\sigma,\gamma_{p}\right)$, see Eq. (57), or simply $U_{e_{1}e_{2}e_{3}}$ for brevity, where the integers $e_{i}$ take values from $1$ to $N_{i}-1$. As discussed earlier, $U_{e_{1}e_{2}e_{3}}$ can also be interpreted as the Wilson operator of an Abelian topological excitation. In this interpretation, the integers $e_{i}$ specify the numbers of $\mathbb{Z}_{N_{i}}$ gauge charges and fluxes carried by the excitation. Equivalently, $U_{e_{1}e_{2}e_{3}}$ can be viewed as the fusion product of $U_{e_{1}}\left(\gamma_{p}\right)$, $U_{e_{2}}\left(\gamma_{p}\right)$, and $U_{e_{3}}\left(\sigma\right)$. Here, $U_{e_{1}}\left(\gamma_{p}\right)$ and $U_{e_{2}}\left(\gamma_{p}\right)$ are $2$-form symmetry operators, while $U_{e_{3}}\left(\sigma\right)$ is a $1$-form symmetry operator. The fusion rules of invertible symmetry operators (with supporting manifolds suppressed) are group-like,
@@
The fusion rules of the non-invertible symmetry operators $L_{m_{1}m_{2}m_{3}}$, for example $L_{ijk}\times L_{rst}$, can be obtained systematically as follows. One first treats the fusion of the unitary parts and the projectors separately. The fusion of the unitary operators is straightforward,
@@
since they appear as multiplicative factors in the functional integral. This result reproduces precisely the unitary component of the symmetry operator $L_{\left(i+r\right)\left(j+s\right)\left(k+t\right)}$. By contrast, the fusion of the projectors yields the product $\mathbb{P}_{ijk}\mathbb{P}_{rst}$, which can be rewritten, inside correlation functions, as a linear combination of invertible symmetry operators, for example as in Eq. (84). Consequently, one can identify a set of invertible operators $U_{abc}$, possibly accompanied by coefficients $f_{abc}$, such that
@@ IV.4.1 $N_{1}=N_{2}=N_{3}=3$, $p=1$
$L_{021}\left(\sigma,\gamma_{p}\right)$ $=\mathcal{U}_{021}\delta\left(2\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\gamma_{p}}a^{1}\right)\delta\left(2\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{2}\right).$ (101)
@@
$\mathbb{P}_{111}=$ $\delta\left(\int_{\gamma_{p}}a^{2}-\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{2}\right).$ (102)
@@ IV.4.2 $N_{1}=4$, $N_{2}=8$, $N_{3}=12$, $N_{123}=4$, $p=1$
The delta functions in $L_{357}$ enforce $\frac{2\pi p}{N_{123}}\left(3k_{2}-5k_{1}\right)=0$, $\frac{2\pi p}{N_{123}}\left(3k_{3}-7k_{1}\right)=0$, and $\frac{2\pi p}{N_{123}}\left(5k_{3}-7k_{2}\right)=0$, all of which are defined modulo $2\pi$. The independent equations are $k_{3}-k_{1}=0\mod 4$ and $k_{2}-3k_{1}=0\mod 4$. Using $4\int_{\gamma_{p}}a^{1}\in 2\pi\mathbb{Z}$, $8\int_{\gamma_{p}}a^{2}\in 2\pi\mathbb{Z}$, and $12\int_{\sigma}b^{3}\in 2\pi\mathbb{Z}$, we can simplified $\mathbb{P}_{357}$ as
@@ IV.4.3 $N_{1}=4$, $N_{2}=8$, $N_{3}=12$, $N_{123}=4$, $p=2$
But now, with $p=2$ and $N_{123}=4$, these delta functions are $\mathbb{P}_{100}\times\mathbb{P}_{100}=\delta\left(4\int_{\gamma_{p}}a^{2}\right)\delta\left(6\int_{\sigma}b^{3}\right)$. These delta functions require that $\int_{\gamma_{p}}a^{2}\in\left\{0,\frac{4\pi}{8},\frac{8\pi}{8},\frac{12\pi}{8}\right\}$ and $\int_{\sigma}b^{3}\in\left\{0,\frac{4\pi}{12},\frac{8\pi}{12},\cdots,\frac{22\pi}{12}\right\}$, and $4\int_{\gamma_{p}}a^{2}\in\left\{0,\pi\right\}$ and $6\int_{\sigma}b^{3}\in\left\{0,\pi\right\}$. Therefore, we can express the delta functions as $\frac{1}{2}\sum_{l_{2}=0}^{1}e^{{\rm i}l_{2}\cdot 4\int_{\gamma_{p}}a^{2}}\times\frac{1}{2}\sum_{l_{3}=0}^{1}e^{{\rm i}l_{3}\cdot 6\int_{\sigma}b^{3}}$. On the other hand, for $L_{200}=\mathcal{U}_{200}\delta\left(\frac{2pN_{2}}{N_{123}}\int_{\gamma_{p}}a^{2}\right)\delta\left(\frac{2pN_{3}}{N_{123}}\int_{\sigma}b^{3}\right)$, the projectors are $\mathbb{P}_{200}=\delta\left(8\int_{\gamma_{p}}a^{2}\right)\delta\left(12\int_{\sigma}b^{3}\right)$, which impose no constraints on the values of $\int_{\gamma_{p}}a^{2}$ and $\int_{\sigma}b^{3}$. This is because $8\int_{\gamma_{p}}a^{2}$ and $12\int_{\sigma}b^{3}$ always equal $0\mod 2\pi$, and the delta functions always return $1$. Therefore,
@@
We summarize the fusion rules for different values of $p$ in Table 2. The fusion rules of invertible symmetry operators are group-like, while those of non-invertible symmetry operators depend on the fusion of the projectors (delta functions). For a general effective field theory $S_{aab}$ with gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$ and a level-$p$ twisted term, the greatest common divisor $N_{123}$ and $p$ determine how a projector can be expressed as a sum of invertible operators in the context of correlation functions, which in turn determines the fusion of the projectors.
@@ V Anomalies of (non-)invertible higher-form symmetries
Below, we first illustrate the mixed anomaly between two higher-form symmetries using the 33d toric code and its effective field theory. We then investigate the mixed anomalies among the generalized higher-form symmetries in the BR topological order described by $S_{aab}$ in Sec. III.
Table 3: Mixed anomalies of generalized symmetries in the effective field theory $S_{aab}$. $J_{e}^{1}=-*a^{1}$, $J_{e}^{2}=-*a^{2}$, and $J_{e}^{3}=*b^{3}$ are conserved currents that generate $2$-, 22-, and $1$-form invertible symmetries, respectively. $J_{m}^{1}=*\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)$, $J_{m}^{2}=*\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right)$, and $J_{m}^{3}=-*\left(a^{3}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\phi^{1}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right)$ are conserved currents that generate $1$-, 11-, and $2$-form non-invertible symmetries, where $d\phi^{1}=\frac{pN_{1}}{N_{123}}a^{1}$, $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$, and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$. To gauge these generalized symmetries, we couple the currents to background fields: 22-form (33-form) background fields are denoted by $\mathcal{B}$ ($\mathcal{C}$). This table lists the outcomes of gauging different pairs of generalized symmetries. The mark $\checkmark$” indicates that there is no mixed anomaly for the corresponding pair. A higher-dimensional bulk term, e.g., $\int_{M_{5}}\frac{1}{2\pi N_{1}}\mathcal{B}_{m}^{1}\mathcal{C}_{e}^{1}$, indicates a mixed anomaly that can be canceled by anomaly inflow from a (4+1)(4+1)D bulk. The mark $\times$” indicates a mixed anomaly manifested by a lack of gauge invariance that cannot be canceled by adding any higher-dimensional bulk term.
@@ V.2 Exploration of mixed anomalies of higher-form symmetries in Borromean-rings topological order
To gauge the $1$-form symmetry generated by $J_{m}^{1}$, we add the coupling term $\mathcal{B}_{m}^{1}\wedge*J_{m}^{1}$ to $S_{aab}$,
@@
$S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{2}\wedge*J_{e}^{2},\quad\begin{cases}\mathcal{C}_{e}^{2}\rightarrow&\mathcal{C}_{e}^{2}+dT_{e}^{2}\\ b^{2}\rightarrow&b^{2}+\frac{1}{N_{2}}T_{e}^{2}\end{cases};$ (139)
$S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\wedge*J_{m}^{2},\quad\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases};$ (140)
$S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{e}^{3}\wedge*J_{e}^{3},\quad\begin{cases}\mathcal{B}_{e}^{3}\rightarrow&\mathcal{B}_{e}^{3}+dV_{e}^{3}\\ a^{3}\rightarrow&a^{3}-\frac{1}{N_{3}}V_{e}^{3}\end{cases};$ (141)
$S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{m}^{3}\wedge*J_{m}^{3},\quad\begin{cases}\mathcal{C}_{m}^{3}\rightarrow&\mathcal{C}_{m}^{3}+dT_{m}^{3}\\ b^{3}\rightarrow&b^{3}+\frac{1}{N_{3}}T_{m}^{3}\end{cases}.$ (142)
We now consider gauging two higher-form symmetries simultaneously by adding two coupling terms to $S_{aab}$. All combinations are summarized in Table 3. Some pairs can be gauged simultaneously and are free of mixed anomalies. In other cases, a mixed anomaly arises because the gauged action fails to remain gauge invariant. Some mixed anomalies can be canceled by coupling to a (4+1)(4+1)D bulk term, while others cannot.
@@
$S_{\text{gauged}}^{e_{1},e_{2}}=$ $S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{1}a^{1}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{2}a^{2},$ (143)
@@
$S_{\text{gauged}}^{e_{1},m_{2}}=$ $S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{1}a^{1}$
@@
$S_{\text{gauged}}^{e_{3},m_{3}}=$ $S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{e}^{3}b^{3}$
@@
$S_{\text{gauging}}^{m_{1},m_{2}}=$ $S_{aab}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)$

### v3

@@ IV.3 An overview of fusion rules of symmetry operators
A general invertible symmetry operator is denoted by $U_{e_{1}e_{2}e_{3}}\left(\sigma,\gamma_{p}\right)$, see Eq. (57), or simply $U_{e_{1}e_{2}e_{3}}$ for brevity, where the integers $e_{i}$ take values from $1$ to $N_{i}-1$. As discussed earlier, $U_{e_{1}e_{2}e_{3}}$ can also be interpreted as the Wilson operator of an Abelian topological excitation. In this interpretation, the integers $e_{i}$ specify the $\mathbb{Z}_{N_{i}}$ gauge charges and fluxes carried by the excitation. Equivalently, $U_{e_{1}e_{2}e_{3}}$ can be viewed as the fusion product of $U_{e_{1}}\left(\gamma_{p}\right)$, $U_{e_{2}}\left(\gamma_{p}\right)$, and $U_{e_{3}}\left(\sigma\right)$. Here, $U_{e_{1}}\left(\gamma_{p}\right)$ and $U_{e_{2}}\left(\gamma_{p}\right)$ are $2$-form symmetry operators, while $U_{e_{3}}\left(\sigma\right)$ is a $1$-form symmetry operator. The fusion rules of invertible symmetry operators (with supporting manifolds suppressed) are group-like,
@@
The fusion rules of the non-invertible symmetry operators $L_{m_{1}m_{2}m_{3}}$, for example $L_{ijk}\times L_{rst}$, can be obtained systematically as follows. We first treat the fusion of the unitary parts and the projectors separately. The fusion of the unitary operators is straightforward,
@@
since they appear as multiplicative factors in the functional integral. This result reproduces precisely the unitary component of the symmetry operator $L_{\left(i+r\right)\left(j+s\right)\left(k+t\right)}$. By contrast, fusing the projectors yields the product $\mathbb{P}_{ijk}\mathbb{P}_{rst}$, which can be rewritten, inside correlation functions, as a linear combination of invertible symmetry operators, for example as in Eq. (84). Consequently, one can identify a set of invertible operators $U_{abc}$, possibly accompanied by coefficients $f_{abc}$, such that
@@ IV.4.1 $N_{1}=N_{2}=N_{3}=3$, $p=1$
$L_{021}\left(\sigma,\gamma_{p}\right)$ $=\mathcal{U}_{021}\delta\left(2\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\gamma_{p}}a^{1}\right)$
$\delta\left(2\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{2}\right).$ (101)
@@
$\mathbb{P}_{111}=$ $\delta\left(\int_{\gamma_{p}}a^{2}-\int_{\gamma_{p}}a^{1}\right)\delta\left(\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{1}\right)$
$\delta\left(\int_{\sigma}b^{3}-\int_{\gamma_{p}}a^{2}\right).$ (102)
@@ IV.4.2 $N_{1}=4$, $N_{2}=8$, $N_{3}=12$, $N_{123}=4$, $p=1$
The delta functions in $L_{357}$ enforce $\frac{2\pi p}{N_{123}}\left(3k_{2}-5k_{1}\right)=0$, $\frac{2\pi p}{N_{123}}\left(3k_{3}-7k_{1}\right)=0$, and $\frac{2\pi p}{N_{123}}\left(5k_{3}-7k_{2}\right)=0$, all of which are defined modulo $2\pi$. The independent equations are $k_{3}-k_{1}=0\mod 4$ and $k_{2}-3k_{1}=0\mod 4$. Using $4\int_{\gamma_{p}}a^{1}\in 2\pi\mathbb{Z}$, $8\int_{\gamma_{p}}a^{2}\in 2\pi\mathbb{Z}$, and $12\int_{\sigma}b^{3}\in 2\pi\mathbb{Z}$, we can simplify $\mathbb{P}_{357}$ as
@@ IV.4.3 $N_{1}=4$, $N_{2}=8$, $N_{3}=12$, $N_{123}=4$, $p=2$
But now, with $p=2$ and $N_{123}=4$, these delta functions are $\mathbb{P}_{100}\times\mathbb{P}_{100}=\delta\left(4\int_{\gamma_{p}}a^{2}\right)\delta\left(6\int_{\sigma}b^{3}\right)$. These delta functions require that $\int_{\gamma_{p}}a^{2}\in\left\{0,\frac{4\pi}{8},\frac{8\pi}{8},\frac{12\pi}{8}\right\}$ and $\int_{\sigma}b^{3}\in\left\{0,\frac{4\pi}{12},\frac{8\pi}{12},\cdots,\frac{22\pi}{12}\right\}$, and $4\int_{\gamma_{p}}a^{2}\in\left\{0,\pi\right\}$ and $6\int_{\sigma}b^{3}\in\left\{0,\pi\right\}$. Therefore, we can express the delta functions as $\frac{1}{2}\sum_{l_{2}=0}^{1}e^{{\rm i}l_{2}\cdot 4\int_{\gamma_{p}}a^{2}}\times\frac{1}{2}\sum_{l_{3}=0}^{1}e^{{\rm i}l_{3}\cdot 6\int_{\sigma}b^{3}}$. On the other hand, for $L_{200}=\mathcal{U}_{200}\delta\left(\frac{2pN_{2}}{N_{123}}\int_{\gamma_{p}}a^{2}\right)\delta\left(\frac{2pN_{3}}{N_{123}}\int_{\sigma}b^{3}\right)$, the projectors are $\mathbb{P}_{200}=\delta\left(8\int_{\gamma_{p}}a^{2}\right)\delta\left(12\int_{\sigma}b^{3}\right)$, which impose no constraints on the values of $\int_{\gamma_{p}}a^{2}$ and $\int_{\sigma}b^{3}$. This is because $8\int_{\gamma_{p}}a^{2}$ and $12\int_{\sigma}b^{3}$ always equal to $0\mod 2\pi$, and the delta functions always return $1$. Therefore,
@@
We summarize the fusion rules for different values of $p$ in Table 2. The fusion rules of invertible symmetry operators are group-like, while those of non-invertible symmetry operators depend on the fusion of the projectors (delta functions). For a general effective field theory $S_{\text{BR}}$ with gauge group $G=\prod_{i=1}^{3}\mathbb{Z}_{N_{i}}$ and a level-$p$ twisted term, the greatest common divisor $N_{123}$ and $p$ determine how a projector can be expressed as a sum of invertible operators in the context of correlation functions, which in turn determines the fusion of the projectors.
@@ V Anomalies of (non-)invertible higher-form symmetries
Below, we first illustrate the mixed anomaly between two higher-form symmetries using the 33d toric code and its effective field theory. We then investigate the mixed anomalies among the generalized higher-form symmetries in the BR topological order described by $S_{\text{BR}}$ in Sec. III.
Table 3: Mixed anomalies of generalized symmetries in the effective field theory $S_{\text{BR}}$. $J_{e}^{1}=-*a^{1}$, $J_{e}^{2}=-*a^{2}$, and $J_{e}^{3}=*b^{3}$ are conserved currents that generate $2$-, 22-, and $1$-form invertible symmetries, respectively. $J_{m}^{1}=*\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)$, $J_{m}^{2}=*\left(b^{2}+\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\beta^{3}-\frac{1}{2}\phi^{1}\frac{N_{3}}{2\pi}b^{3}\right)$, and $J_{m}^{3}=-*\left(a^{3}+\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\phi^{1}-\frac{1}{2}\frac{N_{1}}{2\pi}a^{1}\phi^{2}\right)$ are conserved currents that generate $1$-, 11-, and $2$-form non-invertible symmetries, where $d\phi^{1}=\frac{pN_{1}}{N_{123}}a^{1}$, $d\phi^{2}=\frac{pN_{2}}{N_{123}}a^{2}$, and $d\beta^{3}=\frac{pN_{3}}{N_{123}}b^{3}$. To gauge these generalized symmetries, we couple the currents to background fields: 22-form (33-form) background fields are denoted by $\mathcal{B}$ ($\mathcal{C}$). This table lists the outcomes of gauging different pairs of generalized symmetries. The mark $\checkmark$” indicates that there is no mixed anomaly for the corresponding pair. A higher-dimensional bulk term, e.g., $\int_{M_{5}}\frac{1}{2\pi N_{1}}\mathcal{B}_{m}^{1}\mathcal{C}_{e}^{1}$, indicates a mixed anomaly that can be canceled by anomaly inflow from a (4+1)(4+1)D bulk. The mark $\times$” indicates a mixed anomaly manifested by a lack of gauge invariance that cannot be canceled by adding any higher-dimensional bulk term.
@@ V.2 Exploration of mixed anomalies of higher-form symmetries in Borromean-rings topological order
To gauge the $1$-form symmetry generated by $J_{m}^{1}$, we add the coupling term $\mathcal{B}_{m}^{1}\wedge*J_{m}^{1}$ to $S_{\text{BR}}$,
@@
$S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{2}\wedge*J_{e}^{2},\quad\begin{cases}\mathcal{C}_{e}^{2}\rightarrow&\mathcal{C}_{e}^{2}+dT_{e}^{2}\\ b^{2}\rightarrow&b^{2}+\frac{1}{N_{2}}T_{e}^{2}\end{cases};$ (139)
$S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{2}\wedge*J_{m}^{2},\quad\begin{cases}\mathcal{B}_{m}^{2}\rightarrow&\mathcal{B}_{m}^{2}+dV_{m}^{2}\\ a^{2}\rightarrow&a^{2}-\frac{1}{N_{2}}V_{m}^{2}\end{cases};$ (140)
$S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{e}^{3}\wedge*J_{e}^{3},\quad\begin{cases}\mathcal{B}_{e}^{3}\rightarrow&\mathcal{B}_{e}^{3}+dV_{e}^{3}\\ a^{3}\rightarrow&a^{3}-\frac{1}{N_{3}}V_{e}^{3}\end{cases};$ (141)
$S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{m}^{3}\wedge*J_{m}^{3},\quad\begin{cases}\mathcal{C}_{m}^{3}\rightarrow&\mathcal{C}_{m}^{3}+dT_{m}^{3}\\ b^{3}\rightarrow&b^{3}+\frac{1}{N_{3}}T_{m}^{3}\end{cases}.$ (142)
We now consider gauging two higher-form symmetries simultaneously by adding two coupling terms to $S_{\text{BR}}$. All combinations are summarized in Table 3. Some pairs can be gauged simultaneously and are free of mixed anomalies. In other cases, a mixed anomaly arises because the gauged action fails to remain gauge invariant. Some mixed anomalies can be canceled by coupling to a (4+1)(4+1)D bulk term, while others cannot.
@@
$S_{\text{gauged}}^{e_{1},e_{2}}=$ $S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{1}a^{1}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{2}a^{2},$ (143)
@@
$S_{\text{gauged}}^{e_{1},m_{2}}=$ $S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{C}_{e}^{1}a^{1}$
@@
$S_{\text{gauged}}^{e_{3},m_{3}}=$ $S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{e}^{3}b^{3}$
@@
$S_{\text{gauging}}^{m_{1},m_{2}}=$ $S_{\text{BR}}+\int_{M_{4}}\frac{1}{2\pi}\mathcal{B}_{m}^{1}\left(b^{1}-\frac{1}{2}\frac{N_{2}}{2\pi}a^{2}\beta^{3}+\frac{1}{2}\phi^{2}\frac{N_{3}}{2\pi}b^{3}\right)$